OpenStudy (anonymous):

A triangular based pyramid with base area 35 m^2 and height 8.4 m, calculate the volume. i know that the volume is 1/3 b^2× h

OpenStudy (jdoe0001):

well, then you have both units needed for the volume, plug them in

jimthompson5910 (jim_thompson5910):

Volume of a pyramid V = (1/3)*B*h where V = volume B = area of the base h = height

OpenStudy (anonymous):

Ohh okay I just got confused on the area part but thanks!

OpenStudy (jdoe0001):

1/3 b^2× h b = base, no any of its sides, but the whole base, thus is squared in this case is \(35 m^2\)

OpenStudy (anonymous):

thank you! :))

OpenStudy (jdoe0001):

yw